There are 36 coins in a purse, all nickels and quarters. There are 8 more quarters than nickels. The total amount is $6.20. How many nickels and quarters are there?
Answer by stanbon(26367)

There are 36 coins in a purse, all nickels and quarters. There are 8 more quarters than nickels. The total amount is $6.20. How many nickels and quarters are there?
---------------------
Equations:
Quantity Equation: n + q = 36
Value Equation::::5n +25q = 620 cents
============================================
Multiply thru the Quantity Equation by 5 to get:
5n + 5q = 5*36
---
Subtract that from the Value equation to get:
20q = 620-5*36
20q = 440
q = 22 (number of quarters)
----
Substitute into n+q=36 to solve for "n":
n = 36-22 = 14 (number of nickels)
=====================================
Cheers,
Stan H.